OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure.

OpenStudy (anonymous):

The Given = \[6\sqrt{2}\]

OpenStudy (ivettef365):

the formula is the legs are x and the hyp = xsqrt2

OpenStudy (ivettef365):

so x in this case = 6

OpenStudy (anonymous):

Can you help me with a few more?

OpenStudy (ivettef365):

sure

OpenStudy (anonymous):

Find the hypotenuse of each isosceles right triangle when the legs are of the given measure. Given = \[3\sqrt{2}\]

OpenStudy (ivettef365):

same as the previous, if you have the hyp as 3sqrt2 what is the legs

OpenStudy (anonymous):

6?

OpenStudy (ivettef365):

look at the formula I gave you legs = x hyp = xsqrt2

OpenStudy (ivettef365):

what is the value of x on the hyp of this problem ??

OpenStudy (anonymous):

I honestly do not know v___v Geometry is not my subject

OpenStudy (anonymous):

Like could you work it out and like explain step by step?

OpenStudy (ivettef365):

I am giving you the answer look at it formula legs = x hyp = xsqrt2 hyp = 3sqrt2 (that is what you have as hyp) so now what is the value of x

OpenStudy (anonymous):

But like it says what is the hypotenuse? would I still be looking for the legs?

OpenStudy (ivettef365):

oh ok, now use same formula multiply 3sqrt2 x sqrt2

OpenStudy (anonymous):

Do I need a calculator for that?

OpenStudy (anonymous):

Because if so I don't have one..

OpenStudy (ivettef365):

that = 6, you had the right answer all alone :)

OpenStudy (ivettef365):

do you have multiple choices ?

OpenStudy (anonymous):

No its fill in the blank :/

OpenStudy (anonymous):

and I have 3 more problems like that if you could help...

OpenStudy (ivettef365):

ok

OpenStudy (anonymous):

Find the side of a square whose diagonal is of the given measure. Given =\[15\sqrt{2}\]cm.

OpenStudy (ivettef365):

that is like an isosceles triangle, whose hyp = 15sqrt2, so what is the legs

OpenStudy (anonymous):

So I multiply 15 by the sqrt of 2 right?

OpenStudy (ivettef365):

15 is the side already

OpenStudy (ivettef365):

if they ask you for the are then you multiply

OpenStudy (anonymous):

for the are?

OpenStudy (ivettef365):

??

OpenStudy (anonymous):

I dont know that what you put...

OpenStudy (ivettef365):

the diagonal of a square divides the square in 2 triangles, so the diagonal is the hyp of the triangles, you are asked to find the side, so you have hyp = 15sqrt2 legs = 15 (which are the sides of the square)

OpenStudy (ivettef365):

I which I could draw so I can show you